OpenStudy (anonymous):

Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. Part C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work?

OpenStudy (anonymous):

Can you please help?

OpenStudy (anonymous):

mhm well u just have to add and put words in with your answer that make sence

OpenStudy (anonymous):

It doesnt make sense to me though

OpenStudy (anonymous):

just add the numbers 10 15 and subtrack 40

OpenStudy (anonymous):

But it needs an equation for PArt A

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime.

OpenStudy (anonymous):

just what is aid will solve all of them think bout it

OpenStudy (anonymous):

Maybe 10x + 15w = 40?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Wait a second what i said is wrong

OpenStudy (anonymous):

oh right i forgot something xD

OpenStudy (anonymous):

W = Wages earned, x=amount of hours a week without overtime

OpenStudy (anonymous):

yeah u put them in the wrong order

OpenStudy (anonymous):

10x + 15(40) = W?

OpenStudy (anonymous):

Can someone help please???

OpenStudy (anonymous):

I'll try my best to! Ok so in A we're trying to find the amount of wages earned, W. So we put this at the beginning of the equation. W= Next, we find out that we are working x hours in a week. We already know we make $10 an hour, so add 10x to the equation. W=10x Since there is no overtime, our first equation is finished. Simply W=10x

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What about B it doesn't click in my head, maybe like S=15y + 40? or S=15(40) or what?

OpenStudy (anonymous):

You're very close. Now they tell us we worked 40 hours, so let's include that first. We still make $10 an hour so multiply that by 40. S= 10(40) And here's where you add the 15y S=10(40)+15y.

OpenStudy (anonymous):

And for C, first calculate the maximum amount of money that could be made (without overtime) in a week. To find this, do 10 x 40

OpenStudy (anonymous):

After, subtract that number from 490.

OpenStudy (anonymous):

10 x 40 = 400, 490-400= 90

OpenStudy (anonymous):

Correct! Now take 90 and divide it by 15 to find out how many overtime hours he worked.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

Yep. So he worked 40 hours, plus 6 more overtime hours. This means that Drey worked a total of 46 hours.

OpenStudy (anonymous):

Thank you. Is there a way I can give u a reward or medal?

OpenStudy (anonymous):

Can you also help me with this? negative 1 over 2 x + 2 = –x + 7

OpenStudy (anonymous):

You can click on my profile and click the become a fan button, or click on the best response next to my name for a medal if you wish :)

OpenStudy (anonymous):

And sure

OpenStudy (anonymous):

So like -1/2x + 2 = -x+7

OpenStudy (anonymous):

would i multiply each side by 2/1?

OpenStudy (anonymous):

So like or negative 2/1

OpenStudy (anonymous):

\[\frac{ -1 }{ 2x+2}=-x+7\] Wait, is this what your equation looks like?

OpenStudy (anonymous):

no like this \[-\frac{ 1 }{ 2 }x+2=-x+7\]

OpenStudy (anonymous):

Ah okay. Well first subtract 2 on both sides

OpenStudy (anonymous):

Okay so \[-\frac{ 1 }{ 2 }x=-x + 5?\]

OpenStudy (abhisar):

See here, http://openstudy.com/updates/558469d8e4b091b59af11428

OpenStudy (anonymous):

Yes @nathangroup Now you were correct, multiply both sides by the reciprocal of -1/2, which is 2/1

OpenStudy (anonymous):

So -x=-2x + 10?

OpenStudy (anonymous):

then add 2x to both sides

OpenStudy (anonymous):

x = 10?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Yay!!!! Thank you!

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

I fanned you :D